OpenStudy (anonymous):

find the exact value of sin (-7pi/4)

OpenStudy (anonymous):

does pi stand for 3.14

OpenStudy (anonymous):

-.0958073086

OpenStudy (anonymous):

that should be your awnser

OpenStudy (anonymous):

or should i say ite quals to =-0.096 for your estimate awnser

OpenStudy (anonymous):

Easier Solution: sin(7pi/4) = sin(2pi -- pi/4) = sin(-pi/4) = -sin(pi/4) = - sqrt(2)/2 if you dont have to show your work you could also just look at the unit circle and count over 7 pi/4's and see that its below the x axis. sin of pi/4 = 1/sqrt(2) [some people learn sqrt(2)/2, same thing] and being below the x axis makes it negative.

OpenStudy (phi):

\( - \frac{7 \pi}{4} \) is the same angle if we add \( 2 \pi = \frac{8\pi}{4}\) \[ - \frac{7 \pi}{4} +\frac{8\pi}{4}= \frac{\pi}{4} \] so you want \( \sin\left( \frac{\pi}{4} \right) \) or, in degrees sin(45º)

OpenStudy (phi):

People memorize the sin, cos and tan of 0º, 30º, 45º, 60º and 90º see http://2000clicks.com/MathHelp/GeometryTrigCommonAngles.aspx